### v2

@@ 5.4 Conclusion
[…]
LC is supported by the EPSRC-funded UCL Centre for Doctoral Training in Intelligent, Integrated Imaging in Healthcare (i4health) (EP/S021930/1), and the Wellcome Trust (203147/Z/16/Z and 205103/Z/16/Z). This research was supported by NVIDIA and used NVIDIA RTX A6000 48GB.
Appendix A Physics-based Signal Equations
This appendix details the forward models used to simulate MRI signal intensities from quantitative MRI (qMRI) parameter maps. For each voxel, we assume known proton density (PD), longitudinal relaxation rate ($R_{1}$), and transverse relaxation rate ($R_{2}$ or $R_{2}^{*}$), and optionally a magnetic transfer (MT) parameter. We consider several common MRI sequences, each governed by its own signal equation:
Fast Spin-Echo (FSE).
$S_{\text{FSE}}=B_{1}\cdot\text{PD}\cdot\bigl{(}1-e^{-R_{1}\cdot T_{R}}\bigr{)}\cdot e^{-R_{2}\cdot T_{E}}.$ (1)
Gradient-Echo (GRE).
$S_{\text{GRE}}=B_{1}\cdot\text{PD}\cdot\sin(\alpha)\cdot(1-\text{MT})\cdot\tfrac{1-e^{-R_{1}\cdot T_{R}}}{1-\cos(\alpha)(1-\text{MT})\,e^{-R_{1}\cdot T_{R}}}\cdot e^{-R_{2}^{*}\cdot T_{E}}.$ (2)
Fluid-Attenuated Inversion Recovery (FLAIR).
$S_{\text{FLAIR}}=B_{1}\cdot\text{PD}\cdot e^{-R_{2}\cdot T_{E}}\cdot\bigl{(}1-2\,e^{-R_{1}\cdot T_{I}}+e^{-R_{1}\cdot T_{R}}\bigr{)}.$ (3)
Magnetisation-Prepared Rapid Gradient Echo (MPRAGE).
$S_{\text{MPRAGE}}$ $=\Bigl{|}\sin(\alpha)\,\Bigl{[}\tfrac{1-e^{-R_{1}\cdot T_{R}}}{1-\cos(\alpha)\,e^{-R_{1}\cdot T_{R}}}\,\bigl{(}1-(\cos(\alpha)\,e^{-T_{X}\cdot R_{1}})^{n}\bigr{)}\Bigr{]}\,e^{-T_{D}\cdot R_{1}}$ (4)
$\quad+\,\bigl{(}1-e^{-T_{D}\cdot R_{1}}\bigr{)}\Bigr{|}.$
Here, $B_{1}$ represents the receive field strength, $\alpha$ (in radians) is the flip angle, $T_{R}$ the repetition time, $T_{E}$ the echo time, $T_{I}$ the inversion time, $T_{X}$ the excitation repetition time, $T_{D}$ a delay time, and $n$ the number of excitation pulses (all times in seconds unless noted otherwise).
@@ Appendix A Physics-based Signal Equations
To introduce realistic MRI noise, we add Rician noise Gudbjartsson and Patz (1995) to the simulated signals. Specifically,
$S_{\textrm{noisy}}=\sqrt{(S_{\textrm{MRI}}+\epsilon_{\textrm{r}})^{2}+\epsilon_{\textrm{i}}^{2}}\text{ where }\epsilon_{\textrm{r}}\sim\mathcal{N}(0,\sigma^{2})\text{ and }\epsilon_{\textrm{i}}\sim\mathcal{N}(0,\sigma^{2})$ (5)
where $S_{\text{MRI}}$ is the noise-free signal, and each $\mathcal{N}(0,\sigma^{2})$ term is a zero-mean Gaussian simulating real and imaginary noise components.
All simulations were implemented using the NiTorch libraryhttps://github.com/balbasty/nitorch.
Acquisition Parameters and Data Augmentation
We randomly sampled MRI sequences and parameters from realistic distributions to produce diverse training data, as summarised below. Each sequence was chosen with equal probability to ensure broad coverage of acquisition protocols.
FLAIR:
$T_{E}\sim 10^{\mathcal{N}(\log_{10}(0.02),\log_{10}(0.1))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(5))}\quad\text{s},$
$T_{I}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s}.$
FSE:
$T_{E}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.001),\log_{10}(3))}\quad\text{s}.$
MPRAGE:
$T_{R}\sim\mathcal{N}(23,2.3)\quad\text{s},\quad T_{I}\sim\mathcal{U}(0.6,0.9)\quad\text{s},\quad T_{X}\sim\mathcal{U}(0.004,0.008)\quad\text{s},$
$T_{E}\sim\mathcal{U}(0.002,0.004)\quad\text{s},\quad\alpha\sim\mathcal{U}(5^{\circ},12^{\circ}).$
GRE:
$T_{E}\sim 10^{\mathcal{U}(\log_{10}(0.002),\log_{10}(0.08))}\quad\text{s},\quad T_{R}\sim 10^{\mathcal{U}(\log_{10}(0.005),\log_{10}(5))}\quad\text{s},\quad\alpha\sim\mathcal{U}(5^{\circ},50^{\circ}).$
All sampled parameters were constrained to physically plausible ranges (e.g. taking absolute values as needed). This broad variation in sequences and parameter values helps train the model to handle diverse imaging protocols.
Appendix B Model Architectures
@@
The pre-training setup consists of three components:
  • CNN Encoder:
    • Input: 3D volume with 1 channel
    • 5 downsampling blocks with features: (64, 128, 256, 512, 768)
    • Each block: Two 3D convolutions with kernel size 3
    • Instance normalization and GELU activation after each convolution
    • Dropout rate: 0.2
  • Projector: Two-layer MLP
    • Input dimension: 768
    • Hidden layer (768 → 512) with GELU activation
    • Output projects to contrastive learning space (512 → 128)
  • Reconstructor: Decoder network
    • 4 transposed convolution blocks: (768 → 384 → 192 → 96 → 48)
    • Each upsampling: 2×2×2 stride
    • Final 1×1×1 convolution to output channels

### v3

@@ 5.4 Conclusion
Acknowledgements.
LC is supported by the EPSRC CDT in Intelligent, Integrated Imaging in Healthcare (EP/S021930/1) and by the Wellcome Trust (203147/Z/16/Z and 205103/Z/16/Z). NVIDIA donated the RTX A6000 48 GB GPU used in this research.
Appendix 0.A Physics-based Signal Equations
For each voxel we assume proton density (PD), longitudinal relaxation rate $R_{1}$, transverse relaxation rate $R_{2}$ ($R_{2}^{*}$ for GRE) and, optionally, magnetisation transfer (MT). The receive field is denoted $B_{1}$ and the sequence-specific timing parameters are the repetition time $T_{R}$, echo time $T_{E}$, inversion time $T_{I}$, excitation spacing $T_{X}$, delay $T_{D}$ and excitation count $n$.
Table 4: Forward signal models used for sequence synthesis.
Sequence Signal equation $S=f(\cdot)$
Fast Spin-Echo (FSE) $\text{PD}\,B_{1}\bigl{(}1-e^{-R_{1}T_{R}}\bigr{)}\,e^{-R_{2}T_{E}}$
Gradient-Echo (GRE) $\text{PD}\,B_{1}\sin\!\alpha\,(1-\text{MT})\,\dfrac{1-e^{-R_{1}T_{R}}}{1-\cos\!\alpha\,(1-\text{MT})e^{-R_{1}T_{R}}}\,e^{-R_{2}^{*}T_{E}}$
FLAIR $\text{PD}\,B_{1}e^{-R_{2}T_{E}}\bigl{(}1-2e^{-R_{1}T_{I}}+e^{-R_{1}T_{R}}\bigr{)}$
MPRAGE $\text{PD}\,B_{1}\!\Bigl{|}\sin\!\alpha\frac{1-e^{-R_{1}T_{R}}}{1-\cos\!\alpha\,e^{-R_{1}T_{R}}}\bigl{[}1-(\cos\!\alpha\,e^{-T_{X}R_{1}})^{n}\bigr{]}e^{-T_{D}R_{1}}+1-e^{-T_{D}R_{1}}\Bigr{|}$
@@ Appendix 0.A Physics-based Signal Equations
Rician corruption is applied on-the-fly:
$S_{\text{noisy}}=\sqrt{\bigl{(}S_{\text{MRI}}+\epsilon_{r}\bigr{)}^{2}+\epsilon_{i}^{2}},\quad\epsilon_{r},\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$
as in [13]. All signal synthesis relies on NiTorch [^2] https://github.com/balbasty/nitorch.
Acquisition Parameter Sampling
Table 5: Sampling ranges for synthetic sequence generation ($\log U$ indicates sampling uniformly in log-space).
Sequence $T_{E}$ [s] $T_{R}$ [s] Additional parameters
FLAIR $\log U(0.02,0.10)$ $\log U(0.001,5)$ $T_{I}\sim\log U(0.001,3)$
FSE $\log U(0.001,3)$ $\log U(0.001,3)$
MPRAGE $U(0.002,0.004)$ $N(23,2.3)$ $T_{I}\!\sim\!U(0.6,0.9)$,  $T_{X}\!\sim\!U(0.004,0.008)$,  $\alpha\!\sim\!U(5^{\circ},12^{\circ})$
GRE $\log U(0.002,0.08)$ $\log U(0.005,5)$ $\alpha\!\sim\!U(5^{\circ},50^{\circ})$
All samples are clamped to physically plausible values (negative draws are reflected).
Appendix 0.B Model Architectures
@@
The pre-training setup consists of three components, described in Table 6. NT-Xent projection head is
Table 6: Network modules used during self-supervised pre-training.
Module Layers / blocks Kernel Output dims Notes
CNN encoder 5 conv-blocks $3^{3}$ 64 $\rightarrow$ 768 instance-norm, GELU, dropout 0.2
Projector MLP(768 → 512 → 128) NT-Xent projection head
Reconstructor 4 transposed conv $2^{3}$ 768 $\rightarrow$ 48 L1 reconstruction branch